OpenStudy (anonymous):

( 1/4x ^2 - 2/5 )( - 1/2x - 2/3 ) find the product

OpenStudy (amistre64):

it looks a pain with the fractions

OpenStudy (anonymous):

it is

OpenStudy (amistre64):

( 1/4x ^2 - 2/5 )( - 1/2x - 2/3 ) = N (20)( 1/4x ^2 - 2/5 )( - 1/2x - 2/3 )(6) = (20)(6)N (5x^2- 8)( -3x-4) = 120N ; thats better

OpenStudy (amistre64):

5x^2 - 8 -3x -4 --------- -15x^3 +24x -20x^2 +32 ---------------- \(-15x^3-20x^2+24x+32=120N\) \(\cfrac{-15x^3-20x^2+24x+32}{120}=N\)

OpenStudy (amistre64):

split annd reduce as wanted

OpenStudy (anonymous):

thank u

OpenStudy (amistre64):

youre welcome :)

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=%28+1%2F4+x+^2+++-+2%2F5+%29%28+-+1%2F2+x+++-++2%2F3+%29 if the link lights up, look at the "alternate" forms for the simplification ;)

OpenStudy (anonymous):

ok